2. Let the random variables X and Y have the joint PDF given below: 2e -y 0 xyo0 fxy (x, y) otherwise 0 (a) Find P(X Y < 2) ((c) Find the conditional PDF of Y X x (d) Find P(Y< 3|X = 1)
Answer #1

2) (1) P(x+ J fxy (219) dy che ox Ly=x. (owney T 20x dy die - Bio) ex and ty=2 X+X22 2x=2 x 평뼹 2 (667) *) olva 2-340 (C*) d(6) Marginal pdf of x, L = 12e let dj = 20 2 = 26% 1 co-ex) =26* (€)= 20 22 ,x70 Marginal pdf of y, fy (y) = (a exey de = 20%وے بلام 2xم 2 9 وهم : م ہے کے 29لاک 6 الاحلام - = (/5) رنگ :. و (7) = ((=>/ PC71 ( ومنے لاتے اتے تھے ۔ ہے۔ تھی۔ 2م - - ---- -